# Lattice Boltzmann methods for soft flowing matter

### v1

@@ III.2 Basic ideas of the Lattice Boltzmann method
Let us now introduce the fundamental components that enable the functionality of the LB method. The LB equation states as follows:
@@
The link between lattice Boltzmann equation and Navier-Stokes equation can be found through a Chapman-Enskog analysis Chapman and Cowling (1990), which essentially consists of a multiscale expansion of the distribution function about equilibrium (and of spatial and temporal derivatives) in the Knudsen number $\epsilon=\lambda/L$, where $\lambda$ is the molecular mean-free path and $L$ is a characteristic macroscopic length of the system. For small values of $\epsilon$ ($\epsilon\ll 1$), $\lambda$ is much smaller than $L$, hence a continuum theory would provide a reliable description of the fluid dynamics. It can be shown that this technique allows to recover both continuity and Navier-Stokes equation Krüger et al. (2017); Carenza et al. (2019), in which the kinematic viscosity is given by:
@@
The theory described so far is generally common to different LB methods simulating multiphase and multicomponent fluids at the microscale. However, the details of their implementation, as well as the way in which the physics of the fluid interface (i.e. pressure and surface tension) is computationally modeled, depends significantly on the type of LB adopted. In the next paragraph, we describe three widely used LB approaches to simulate soft flowing systems, namely color-gradient, free energy, and pseudopotential models.
@@ III.3 Modern Lattice Boltzmann methods for soft matter
III.3.1 Color-gradient LB
@@
where $f_{i}^{*}=\sum_{k}f_{i}^{k,*}$ denotes the set of post-perturbation distributions, $\rho=\rho_{A}+\rho_{B}$, $\theta_{i}$ is the angle between the phase field gradient and the $i^{th}$ lattice vector, $f_{i}^{eq,0}=\sum_{k}f_{i}^{k}(\rho,{\bf u}=0)^{eq}$ is the total zero-velocity equilibrium distribution function Leclaire et al. (2012) and $\beta$ is a free parameter tuning the interface width, thus playing the role of an inverse diffusion length scale Latva-Kokko and Rothman (2005). In Montessori et al. (2018a), it has also been shown that the color gradient LB scheme can be further stabilized by filtering out the high-order non-hydrodynamic (ghost) modes emerging in the under-relaxed regime (i.e. $\tau\geq\Delta t$) Montessori et al. (2015); Zhang et al. (2006); Latt and Chopard (2006); Coreixas et al. (2017); Mattila et al. (2017); Hegele et al. (2018); Montessori et al. (2016). This improvement essentially allows the color gradient to operate at higher values of droplet speed minimizing the effects due to spurious currents and ghost modes Succi et al. (1995); Benzi et al. (1992); Montessori et al. (2018a, b).
Near-contact interactions.
@@
By neglecting surface tension variations along the interface (thus the stress tensor ${\bf T}$ can be approximated as $\mathbf{T}=-p\mathbf{I}$ (Brackbill et al. (1992)) and by projecting Eq.(40) along the normal at the interface, an extended Young-Laplace equation can be obtained (Chan et al. (2011); Williams and Davis (1992))
@@
We finally mention that, over the years, other methods have been proposed to model the physics of near-contact forces in these systems Nekovee et al. (2000); Love et al. (2001); Chen et al. (2000). However, unlike the approach discussed here where their effect is taken into account by introducing a repulsive force (as a mesoscale representation of supramolecular interactions), in these alternative methods two additional BGK-like equations are used, one for the distribution functions of a surfactant (which essentially prevents droplet coalescence) and another one for the relaxation of the average dipole vector. The latter is because the interaction among surfactant molecules depends on their relative distance and on their dipolar orientation. An important difference between these two approaches is that the color-gradient method augmented with near-contact forces only requires two sets of distribution functions regardless of the number of droplets, thus allowing for the simulations of mixtures with high volume fractions of droplets (see Section V) at a dramatically reduced computational cost.
@@ III.3.2 Free-energy model
Alongside the continuum equations, in the multicomponent free energy model one also needs an evolution equation for the phase field $\phi$. This is governed by the Cahn-Hilliard equation Cahn and Hilliard (1958)
@@
Multiphase field model.
In the same spirit of the color-gradient LB approach, the free-energy model can also be extended to simulate systems with more than two components and where coalescence is inhibited, such as dense emulsions Foglino et al. (2017). An emulsion of $N$ immiscible droplets, for example, can be described as a set of scalar fields $\phi_{n}$ ($n=1,....,N$) capturing the density of each droplet, whose dynamics obeys a corresponding set of advection-diffusion equation
@@
where the first term guarantees the existence of the minima $\phi_{n}=\phi_{0}$ (for example inside the $n$-th droplet) and $\phi_{n}=0$ (outside), and the second one accounts for the fluid interface. The positive constants $a$ and $k$ control surface tension and interface thickness, given by Eq.(46) and Eq.(49), respectively. Finally, the last term models a repulsive effect hindering droplet merging. This contribution slightly modifies the stress tensor, which includes an additional term proportional to $-\sum_{n}\phi_{n}\nabla\mu_{n}$.
Extension to active fluids.
In recent years, research on active matter in general and active fluids in particular has known a burgeoning growth Marchetti et al. (2013). Active matter presents a few distinctive features that set it apart from passive one, such as lack of reciprocity of pair interactions and the resulting breaking of time-reversal symmetry which triggers spontaneous motion. More often than not, however, the actual physical mechanisms responsible for the above features can be described in terms of additional internal degrees of freedom, coupled to the external ones (position and momentum) describing passive matter. As a result, from an operational standpoint, active systems can be regarded as passive systems with internal degrees of freedom coupled to the external ones in such a way as to sustain feedback loops with the environment. Therefore, the lattice Boltzmann method can be readily extended to handle active flowing matter as well. Indeed, over the last two decades, the free-energy lattice Boltzmann method has been successfully adapted to the simulation the hydrodynamics of active fluids and self-propelled droplets, in which the active fluid is segregated Carenza et al. (2019); Cates et al. (2009). As compared to an isotropic passive fluid, active fluids are made of individually oriented units supporting spontaneous symmetry breaking of rotational order, i.e. capable of developing orientationally ordered structures on macroscopic scales. This feature is generally captured by introducing an order parameter, such as the polar vector ${\bf p}({\bf x},t)$ or the tensor ${\bf Q}({\bf x},t)$, describing the ordering properties of the symmetry-broken phase, such as liquid crystals de Gennes and Prost (1993).
@@
Basic implementation.
Historically, two main free-energy lattice Boltzmann schemes have been used to simulate binary fluids. In the first one, the Cahn-Hilliard equation is solved by introducing a set of distribution functions $g_{i}({\bf x},t)$ connected to the phase field and the Navier-Stokes equation via the usual distribution functions $f_{i}({\bf x},t)$ defining zeroth, first and second order momenta. In the second one, the approach is hybrid, meaning that the Cahn-Hilliard equation is integrated via finite difference algorithms, while the Navier-Stokes is through a standard LB method. Both approaches have been widely used to simulate a variety of soft matter systems under confinement, such as liquid-vapor mixtures Gonnella et al. (2007), binary and ternary fluids Lamura et al. (1999); Xu et al. (2003, 2006); Gonnella et al. (2010); Tiribocchi et al. (2011a, 2021a, 2020a, 2021b, 2021c), liquid crystals Cates et al. (2009); Stratford et al. (2014); Lesniewska et al. (2024); Denniston et al. (2004); Sulaiman et al. (2006); Marenduzzo et al. (2004, 2008); Tiribocchi et al. (2011b, 2010, 2014, 2013) and active matter Marenduzzo et al. (2007a); Cates et al. (2008); Shendruk et al. (2017); Marenduzzo et al. (2007b); De Magistris et al. (2014); Negro et al. (2019); Bonelli et al. (2016, 2019); Tiribocchi et al. (2023a, b). Since some applications discussed in Section V are simulated using the second algorithm, we shortly recap its main features, while a detailed description of the first one can be found, for example, in Refs.Krüger et al. (2017); Carenza et al. (2019).
@@
    where $\Delta t_{FD}$ is the time step of the finite difference scheme and the velocity ${\bf u}$ stems from the lattice Boltzmann equation. Note that the terms on the right-hand side are computed at time $t$ and lattice position ${\bf x}$.
@@ III.3.3 Lattice pseudo-potential approach
${\bf F}_{\sigma}({\bf x})=-\psi_{\sigma}({\bf x})\sum_{\tilde{\sigma}}G_{\sigma\tilde{\sigma}}\sum_{i}w_{i}\psi_{\tilde{\sigma}}({\bf x}+{\bf c}_{i}\Delta t){\bf c}_{i}\Delta t,$ (63)
where the sum runs over the lattice links, $\sigma$ and $\tilde{\sigma}$ are the two interacting components and $G_{\sigma\tilde{\sigma}}$ is a parameter setting the strength of their interaction and controlling the surface tension in the model. Unlike the free energy LB, here the surface tension is an emergent effect and can be explicitly calculated using the Young-Laplace test Krüger et al. (2017).
Thermodynamic consistency.
@@ A Taylor expansion of $\psi({\bf x}+{\bf c}_{i}\Delta t)$ gives
$\psi({\bf x}+{\bf c}_{i}\Delta t)=$ $\psi({\bf x})+c_{i\alpha}\Delta t\partial_{\alpha}\psi({\bf x})$ (65)
@@
$+$ $\frac{1}{6}c_{i\alpha}c_{i\beta}c_{i\gamma}\Delta t^{3}\partial_{\alpha}\partial_{\beta}\partial_{\gamma}\psi({\bf x})+...$
Substituting Eq.(65) into Eq.(64) leads to
$F_{\alpha}({\bf x})=$displaystyle-- $G\psi({\bf x})\sum_{i}w_{i}c_{i\alpha}\Delta t\biggl{(}\psi({\bf x})+c_{i\alpha}\Delta t\partial_{\alpha}\psi({\bf x})$ (66)
$+$ $\frac{1}{2}c_{i\alpha}c_{i\beta}\Delta t^{2}\partial_{\alpha}\partial_{\beta}\psi({\bf x})+....\biggr{)}.$
@@
$F_{\alpha}({\bf x})=-G\psi({\bf x})\left(c_{s}^{2}\Delta t^{2}\nabla\psi({\bf x})+\frac{c_{s}^{4}\Delta t^{4}}{2}\nabla\nabla^{2}\psi({\bf x})\right).$ (67)
@@
where $p_{EOS}(\rho)$ is the desired equation of state. Such an approach, combined with a recent version of the entropic LB Hosseini et al. (2023) and a higher-order isotropic discretization of the gradient of the pseudo-potential, has permitted to simulate multiphase flows at Weber and Reynolds numbers above $10^{2}$ and $10^{3}$ respectively Montessori et al. (2017), unattainable by standard pseudo-potential methods.
@@ III.4 Boundary conditions
We conclude this section with a short discussion on the boundary conditions used in the application (see Section V). Since the systems under study are either unbounded or are confined within a microchannel geometry, we focus on periodic boundaries and on solid ones modeling parallel flat walls, while we refer to specific textbooks Succi (2018); Krüger et al. (2017) for a general overview. It is worth noting that, despite the complexity of the systems under study, often involving confined deformable interfaces in close contact, the implementation of the boundary conditions (at least for the two aforementioned cases) follows well-known standard procedures. This is actually a key strength of the method, which proves capable of capturing the mesoscale physics of different examples of confined soft matter with minimal modifications to the existing algorithms.
@@
Figure 9: The figure is inspired by Ref.Succi (2018); Krüger et al. (2017).
@@
$\rho u_{x,w}$ $=$ $\sum_{i}f_{i}c_{ix}=0\rightarrow f_{8}(t)-f_{7}(t)=$
$\hskip 69.70915ptf_{6}(t)-f_{5}(t)+f_{3}(t)-f_{1}(t),$
$\rho u_{y,w}$ $=$ $\sum_{i}f_{i}c_{iy}=0\rightarrow f_{7}(t)+f_{4}(t)+f_{8}(t)=$
$\hskip 69.70915ptf_{5}(t)+f_{2}(t)+f_{6}(t).$
@@
Further extensions of this approach aimed at guaranteeing improved numerical stability have also been proposed Lamura and Gonnella (2001); Tiribocchi et al. (2011a).

### v2

@@ III.2 Basic ideas of the Lattice Boltzmann method
Let us now introduce the fundamental components that enable the functionality of the LB method. The LB equation states as follows
@@
The link between LB equation and Navier-Stokes equation can be found through a Chapman-Enskog analysis Chapman and Cowling (1990), which essentially consists of a multiscale expansion of the distribution function about equilibrium (and of spatial and temporal derivatives) in the Knudsen number $\epsilon=\lambda/L$, where $\lambda$ is the molecular mean-free path and $L$ is a characteristic macroscopic length of the system. For small values of $\epsilon$ ($\epsilon\ll 1$), $\lambda$ is much smaller than $L$, hence a continuum theory would provide a reliable description of the fluid dynamics. It can be shown that this technique allows to recover both continuity and Navier-Stokes equation Krüger et al. (2017); Carenza et al. (2019), in which the kinematic viscosity is given by
@@
The theory described so far is generally common to different LB methods simulating multiphase and multicomponent fluids at the microscale. However, the details of their implementation, as well as the way in which the physics of the fluid interface (i.e. pressure and surface tension) is computationally modeled, depends significantly on the type of LB adopted. In the next paragraph, we describe three widely used LB approaches to simulate soft-flowing systems, namely color-gradient, free energy, and pseudopotential models.
@@ III.3 Modern Lattice Boltzmann methods for soft matter
III.3.1 Color-gradient Lattice Boltzmann
@@
where $f_{i}^{*}=\sum_{k}f_{i}^{k,*}$ denotes the set of post-perturbation distributions, $\rho=\rho_{A}+\rho_{B}$, $\theta_{i}$ is the angle between the phase field gradient and the $i^{th}$ lattice vector, $f_{i}^{eq,0}=\sum_{k}f_{i}^{k}(\rho,{\bf u}=0)^{eq}$ is the total zero-velocity equilibrium distribution function Leclaire et al. (2012) and $\beta$ is a free parameter tuning the interface width, thus playing the role of an inverse diffusion length scale Latva-Kokko and Rothman (2005). In Montessori et al. (2018a), it has been also shown that the color gradient LB scheme can be further stabilized by filtering out the high-order non-hydrodynamic (ghost) modes emerging in the under-relaxed regime (i.e. $\tau\geq\Delta t$) Montessori et al. (2015); Zhang et al. (2006); Latt and Chopard (2006); Coreixas et al. (2017); Mattila et al. (2017); Hegele et al. (2018); Montessori et al. (2016). This improvement essentially allows the model to operate at higher values of droplet speed minimizing the effects due to spurious currents and ghost modes Succi et al. (1995); Benzi et al. (1992); Montessori et al. (2018a, b).
Near-contact interactions
@@
By neglecting surface tension variations along the interface (thus the stress tensor ${\bf T}$ can be approximated as $\mathbf{T}=-p\mathbf{I}$ Brackbill et al. (1992) and by projecting Eq.(40) along the normal at the interface, an extended Young-Laplace equation can be obtained Chan et al. (2011); Williams and Davis (1992)
@@
We finally mention that, over the years, other methods have been proposed to model the physics of near-contact forces in these systems Nekovee et al. (2000); Love et al. (2001); Chen et al. (2000). However, unlike the approach discussed here where their effect is taken into account by introducing a repulsive force (as a mesoscale representation of supramolecular interactions), in these alternative methods two additional BGK-like equations are used, one for the distribution functions of a surfactant (which essentially prevents droplet coalescence) and another one for the relaxation of the average dipole vector. The latter is because the interaction among surfactant molecules depends on their relative distance and on their dipolar orientation. An important difference between these two approaches is that the color-gradient method augmented with near-contact forces only requires two sets of distribution functions regardless of the number of droplets, thus allowing for the simulations of mixtures with high droplet volume fractions (see Section V) at a dramatically reduced computational cost.
@@ III.3.2 Free-energy model
Alongside the continuum equations, in the multicomponent free energy model one also needs an evolution equation for the phase field $\phi$. This one is governed by the Cahn-Hilliard equation Cahn and Hilliard (1958)
@@
Multiphase field model
In the same spirit of the color-gradient LB approach, the free-energy model can also be extended to simulate systems, such as dense emulsions Foglino et al. (2017), with more than two components and where coalescence is inhibited. An emulsion of $N$ immiscible droplets, for example, can be described as a set of scalar fields $\phi_{n}$ ($n=1,....,N$) capturing the density of each droplet, whose dynamics obeys a corresponding set of advection-diffusion equation
@@
where the first term guarantees the existence of the minima $\phi_{n}=\phi_{0}$ (for example inside the $n$-th droplet) and $\phi_{n}=0$ (outside), and the second one accounts for the fluid interface. The positive constants $a$ and $k$ control surface tension and interface thickness, given by Eq.(46) and Eq.(49), respectively. Finally, the last term models a repulsive effect hindering droplet merging. This contribution slightly modifies the stress tensor of Eq.LABEL:Nav_Stok_eq, which includes a term proportional to $-\sum_{n}\phi_{n}\nabla\mu_{n}$.
Extension to active fluids
In recent years, research on active matter in general and active fluids in particular has known a burgeoning growth Marchetti et al. (2013). Active matter presents a few distinctive features that set it apart from passive one, such as lack of reciprocity of pair interactions and the resulting breaking of time-reversal symmetry which triggers spontaneous motion. More often than not, however, the actual physical mechanisms responsible for the above features can be described in terms of additional internal degrees of freedom, coupled to the external ones (position and momentum) describing passive matter. As a result, from an operational standpoint, active systems can be regarded as passive systems with internal degrees of freedom coupled to the external ones in such a way as to sustain feedback loops with the environment. Therefore, the LB method can be readily extended to handle active flowing matter as well. Indeed, over the last two decades, the free-energy LB method has been successfully adapted to the simulation the hydrodynamics of active fluids and self-propelled droplets, in which the active fluid is segregated Carenza et al. (2019); Cates et al. (2009). As compared to an isotropic passive fluid, active fluids are made of individually oriented units supporting spontaneous symmetry breaking of rotational order, i.e. capable of developing orientationally ordered structures on macroscopic scales. This feature is generally captured by introducing an order parameter, such as the polar vector ${\bf p}({\bf x},t)$ or the tensor ${\bf Q}({\bf x},t)$, describing the ordering properties of the symmetry-broken phase, such as liquid crystals de Gennes and Prost (1993).
@@
Basic implementation
Historically, two main free-energy LB schemes have been used to simulate binary fluids. In the first one, the Cahn-Hilliard equation is solved by introducing a set of distribution functions $g_{i}({\bf x},t)$ connected to the phase field and the Navier-Stokes equation via the usual distribution functions $f_{i}({\bf x},t)$ defining zeroth, first and second order momenta. In the second one the approach is hybrid, meaning that the Cahn-Hilliard equation is integrated via finite difference algorithms, while the Navier-Stokes one through a standard LB method. Both approaches have been widely used to simulate a variety of soft matter systems under confinement, such as liquid-vapor mixtures Gonnella et al. (2007), binary and ternary fluids Lamura et al. (1999); Xu et al. (2003, 2006); Gonnella et al. (2010); Tiribocchi et al. (2011a, 2021a, 2020a, 2021b, 2021c), liquid crystals Cates et al. (2009); Stratford et al. (2014); Lesniewska et al. (2024); Denniston et al. (2004); Sulaiman et al. (2006); Marenduzzo et al. (2004, 2008); Tiribocchi et al. (2011b, 2010, 2014, 2013) and active matter Marenduzzo et al. (2007a); Cates et al. (2008); Shendruk et al. (2017); Marenduzzo et al. (2007b); De Magistris et al. (2014); Negro et al. (2019); Bonelli et al. (2016, 2019); Tiribocchi et al. (2023a, b). Since some applications discussed in Section V are simulated using the second algorithm, we shortly recap its main features, while a detailed description of the first one can be found, for example, in Refs.Krüger et al. (2017); Carenza et al. (2019).
@@
    where $\Delta t_{FD}$ is the time step of the finite difference scheme and the velocity ${\bf u}$ stems from the LB equation. Note that the terms on the right-hand side are computed at time $t$ and lattice position ${\bf x}$.
@@ III.3.3 Lattice pseudo-potential approach
${\bf F}_{k}({\bf x})=-\psi_{k}({\bf x})\sum_{\tilde{k}}G_{k\tilde{k}}\sum_{i}w_{i}\psi_{\tilde{k}}({\bf x}+{\bf c}_{i}\Delta t){\bf c}_{i}\Delta t,$ (63)
where the sum runs over the lattice links, $k$ and $\tilde{k}$ are the two interacting components and $G_{k\tilde{k}}$ is a parameter setting the strength of their interaction and controlling the surface tension in the model. Unlike the free energy LB, here the surface tension is an emergent effect and can be explicitly calculated using the Young-Laplace test Krüger et al. (2017).
Thermodynamic consistency
@@ A Taylor expansion of $\psi({\bf x}+{\bf c}_{i}\Delta t)$ gives
$\psi({\bf x}+{\bf c}_{i}\Delta t)=\psi({\bf x})$ $+$ $c_{i\alpha}\Delta t\partial_{\alpha}\psi({\bf x})$
@@
$+$ $\frac{1}{6}c_{i\alpha}c_{i\beta}c_{i\gamma}\Delta t^{3}\partial_{\alpha}\partial_{\beta}\partial_{\gamma}\psi({\bf x})+....$
Substituting Eq.(III.3.3) into Eq.(64) leads to
${\bf F}({\bf x})=$displaystyle-- $G\psi({\bf x})\sum_{i}w_{i}{\bf c}_{i}\Delta t\biggr{(}\psi({\bf x})+c_{i\alpha}\Delta t\partial_{\alpha}\psi({\bf x})$ (66)
$+$ $\frac{1}{2}c_{i\alpha}c_{i\beta}\Delta t^{2}\partial_{\alpha}\partial_{\beta}\psi({\bf x})+....\biggl{)}.$
@@
${\bf F}({\bf x})=-G\psi({\bf x})\left(c_{s}^{2}\Delta t^{2}\nabla\psi({\bf x})+\frac{c_{s}^{4}\Delta t^{4}}{2}\nabla\nabla^{2}\psi({\bf x})\right).$ (67)
@@
where $p_{EOS}(\rho)$ is the desired equation of state. Such an approach, combined with a recent version of the entropic LB Hosseini et al. (2023) and a higher-order isotropic discretization of the gradient of the pseudo-potential, allows for the simulation of multiphase flows at Weber and Reynolds numbers above $10^{2}$ and $10^{3}$ respectively Montessori et al. (2017), unattainable by standard pseudo-potential methods.
@@ III.4 Boundary conditions
We conclude this section with a short discussion on the boundary conditions used in the applications presented in Section V. Since the systems under study are either unbounded or confined within a microchannel, we focus on periodic boundaries and on solid ones modeling parallel flat walls, while we refer to specific textbooks Succi (2018); Krüger et al. (2017) for a general overview. It is worth noting that, despite the complexity of the systems under study often involving confined deformable interfaces in close contact, the implementation of the boundary conditions (at least for the two aforementioned cases) follows well-known standard procedures. This is actually a key strength of the method, which proves capable of capturing the mesoscale physics of different examples of confined soft matter with minimal modifications of existing algorithms.
@@
Figure 9: In this figure, dotted spots represent solid nodes located above the top wall. Red dotted arrows indicate unknown distribution functions. The figure is inspired by Ref.Succi (2018); Krüger et al. (2017).
@@
$\rho u_{x,w}$ $=$ $\sum_{i}f_{i}c_{ix}=0\rightarrow$
$f_{8}(t)-f_{7}(t)=f_{6}(t)-f_{5}(t)+f_{3}(t)-f_{1}(t),$
$\rho u_{y,w}$ $=$ $\sum_{i}f_{i}c_{iy}=0\rightarrow$
$f_{7}(t)+f_{4}(t)+f_{8}(t)=f_{5}(t)+f_{2}(t)+f_{6}(t).$
@@
Further extensions of this approach aimed at guaranteeing improved numerical stability have been also proposed Lamura and Gonnella (2001); Tiribocchi et al. (2011a).